Jungkook invested a certain amount of money at a 10% interest rate and Php 2000 more than that at a 12% interest rate. His total annual interest earned is Php 1340. How much money did he invest at 12% interest rate?


1
Expert's answer
2021-05-07T11:29:40-0400

Jungkook invested a certain amount of money at 10%

interest rate and $2000 more than that at 12% interest rate.

her total yearly interest eanrd is $1340.how much did she invest at each rate?


let   x amount invested at 10%


 then  x+2000 invested at 12%


[interest earned 10%]+[interest earned 12%]=1340


         

        (x)10% + (x+2000) 12% = 1340

            

      "(x)(0.10) +(x+2000)(0.12)= 1340"


       "0.1x+0.12x+240=1340"

       "0.22x=1100"


         "x=\dfrac{1100}{0.22}=5000"


The amount of money he invests at 12% is (5000+2000)=$7000